# PsycoLLM: Enhancing LLM for Psychological Understanding and Evaluation

### v2

@@ III-D Supervised Fine-Tuning
The hyperparameters in this process are configured as follows: the learning rate is set to 1e-5, the batch size is set to 128, and the number of epochs is set to 3. We employ full parameter fine-tuning to train the model using eight A6000 GPUs, each equipped with 48GB memory.
@@ IV Benchmark
Figure 5: Examples of MCQs in the proposed benchmark.
@@ IV-A Dataset Collection and Preprocessing
In order to accurately evaluate the model performance in the psychology domain, we introduce a benchmark that psychological professional individuals need to master. The data is sourced from publicly available examination questions. We follow a standardized data preprocessing procedure. For some data, we first need to perform optical character recognition (OCR) to convert images to text. After that, we invite several students to manually review the collected data to ensure its quality and consistency with the original document. This process involves rectifying formatting errors, eliminating duplicate questions, and rectifying any instances of garbled characters. Our proposed psychological benchmark draws inspiration from the format of the most authoritative psychological counseling examination in China, comprising two primary components: objective questions and subjective questions. Note that the objective questions mainly comprise two types of multiple-choice questions (MCQs): MCQ with only a single correct option (SMCQ) and MCQ with multiple correct options (MMCQ). The content of benchmark can be categorized into three sections: professional ethics, theoretical proficiency, and case analysis. The professional ethics section generally encompasses general principles such as honesty, integrity, professional responsibility, and social responsibility. Theoretical proficiency typically pertains to fundamental theories and concepts in psychology, including cognitive psychology, developmental psychology, and related fields. Both the professional ethics and theoretical proficiency sections contain only MCQs, designed to evaluate the mastery of knowledge within the psychological domain. For case analysis, it includes both MCQs and iterative questioning based on given case, aiming to better evaluate performance in real-world scenarios. The former focuses on assessing the grasp of professional knowledge related to the case, while the latter offers a more open-ended assessment of the model’s capability in practical utility. Finally, we obtained a total of 3863 MCQs and 20 cases with 100 QAs. An example of MCQs is shown in Fig. 5, and an example of case-based QA is presented in Fig. 6. The statistics of the benchmark are shown in Table III, detailing the distribution of questions across various categories. The questions are classified into three main categories: Level 2 Psychological Counselor Exam, Level 3 Psychological Counselor Exam, and Others. Note that, according to the official definitions, the Level 2 exam is considered more challenging than the Level 3 exam.
@@ IV-B Evaluation Metric
There are three elements in each instance of the case analysis. $\mathcal{B}$ is the description of the given case, including the background, environment, etc. of the seeker, $\mathcal{Q}$ is the question according to the case, such as the diagnosis, reason and treatment for the case and $\mathcal{A}$ is the reference answer for each questions. For QA in case analysis, we use the models to generate the response $\mathcal{O}$ according to the given $\mathcal{B}$ and $\mathcal{Q}$. We evaluate the response $\mathcal{O}$ from model with the corresponding references $\mathcal{A}$. We follow existing text generation metric, including ROUGE-1 (R-1), ROUGE-L (R-L) [53], BLEU-4 (B-4) [54], BERTScore [55]. For the ROUGE metric, we report the $\textsc{F}_{1}$ scores for both R-1 and R-L. R-1 measures unigram overlap, which serves as an indicator of informativeness, while R-L evaluates the longest common subsequence overlap, providing an assessment of fluency. BERTScore leverages embedding from pre-trained encoder to capture the semantic and contextual information of tokens through computing cosine similarities. This metric allows for a more nuanced evaluation of text generation quality by considering the underlying meaning and context rather than just surface-level token matching.
For SMCQ, we use the standard accuracy to evaluate the model performance. For MMCQ, we employ two types of accuracy metrics: standard accuracy and elastic accuracy. Standard accuracy is defined as the ratio of completely correct answers to the total number of questions, providing a clear measure of the model’s precision by counting only fully accurate responses. Elastic accuracy, on the other hand, accounts for partially correct answers that contain no incorrect options but are not entirely complete. This metric offers a evaluation by recognizing the model’s ability to produce answers that are mostly correct, even if they are not fully comprehensive. The elastic accuracy for a MMCQ is computed as follows:
@@ V-A Baseline
  • BaiChuan2555https://github.com/baichuan-inc/Baichuan2: This is a generation of open-source large language models launched by Baichuan Intelligent Technology, which is trained on a high-quality corpus with 2.6 trillion tokens.
@@
  • LLaMA3-Chinese [56] [^6] https://github.com/seanzhang-zhichen/llama3-chinese: This model uses LLama3 as the base model and is fine-tuned with high-quality Chinese instruction datasets with the training methods of DORA and LORA+.
  • Yi (e.g., Yi-1.5-6B, Yi-1.5-9B) [^7] https://github.com/01-ai/Yi: These open-source large language models are trained from scratch by developers at 01.AI, which are trained with a high-quality corpus of 500B tokens.
@@
Besides, we also compare with several psychological LLMs:
@@ V-B Evaluation Details
Prompt in our experiments usually comprises following components: a well-defined task description and a pertinent question soliciting the model’s response. For tasks involving MCQ, we extract answers from model outputs using an empirically designed regular expression.
@@ VI-B Results on Case-Based QA
We report the results of subjective questions in Table V. The results obtained from automatic evaluations using ROUGE and BLEU metrics for these models are relatively low. This may be due to the inherent limitations of these metrics in capturing the nuances and complexities of natural language generation, as they primarily focus on surface-level n-gram overlap rather than semantic equivalence and contextual appropriateness. Therefore, we also incorporate BERTScore into our evaluation framework. Most models consistently score between 63 and 65 on the BS indicator, indicating the consistency and stability of the BS evaluation. PsycoLLM demonstrates superior performance on R-1, B-4, and BERTScore, respectively and EmoLLM achieves better results on R-L. Additionally, although EmoLLM and MindChat do not perform well on MCQs, they achieve relatively strong performance in R-1, R-L, and B-4. The main reason might be that these models are fine-tuned with psychological conversation data, enabling them to better grasp case information and generate higher-quality responses.
We provide a comparative example of the results from PsycoLLM and GPT-4o in Fig. 6. Both models are proficient at identifying the majority of patient issues. A notable distinction emerges in their responses: GPT-4o tends to offer more detailed and specific answers, while PsycoLLM generally provides more concise responses.
@@ VI-C Effect of Difficulty
VI-D Discussion
Model Size
On the one hand, within the same model families, models with larger parameters generally obtain better results in Table IV. For instance, models such as YI-1.5-9B and Qwen1.5-14B-Chat demonstrate better performance compared to their smaller counterparts, YI-1.5-6B and Qwen1.5-7B-Chat, respectively. This performance enhancement can be attributed to the increased capacity of models with larger parameters to capture complex patterns and effectively learn text representations from the data. This observation aligns with findings from a broad spectrum of existing studies that emphasize the advantages of larger model sizes in terms of learning capacity and performance. On the other hand, when compared models from different model families, this findings have slight difference. For example, YI-1.5-9B outperform Baichuan2-13B and Qwen1.5-14B-Chat in the MCQ evaluation. This discrepancy shows that training corpus, and other factors such as training procedures also significantly influence performance.
Model preference in MMCQs
In our experiments, we observe distinct response patterns from different models when addressing MMCQs. For example, GPT-3.5-turbo and GPT-4o frequently provide a single option in MMCQs, whereas Qwen-turbo tends to select all available options. Therefore, GPT-3.5-turbo and GPT-4o have a large improvements when shifting from standard accuracy measure to elastic accuracy metric. The main reason might be that, although GPT-3.5-turbo and GPT-4 are multilingual, they primarily focus on English. In English-speaking contexts, MMCQs are less prevalent compared to SMCQs. Consequently, the training data for these models may predominantly contains SMCQs and lack a substantial MMCQ dataset. This discrepancy leads the models to exhibit a tendency to select only one answer when presented with MMCQs.
Fine-tuned Models vs. General Models
Among these models, EmoLLM, MindChat are fine-tuned with the psychology-related dataset and others are for general domain. Specifically, EmoLLM and MindChat leverage empathetic conversation data for fine-tuning, which contributes to their relatively higher performance in Case-Based QA than general LLMs. However, due to their limited incorporation of knowledge-based data, their performance in MCQ evaluations is average.
@@ VII Conclusion
In this paper, to enhance LLM in the psychological domain, we propose a high-quality psychological dataset. Specifically for multi-turn dialogue, we employ a comprehensive pipeline that includes stages of generation, evidence support, and refinement. For knowledge-based QA, we use a teacher-student LLM-based agent to enhance data generation. Additionally, we develop a benchmark based on psychological examinations. Fine-tuning LLMs with the proposed dataset results in the development of PsycoLLM, which demonstrates superior performance on the benchmark, excelling in both knowledge mastery and case analysis compared to other LLMs.
This study primarily focuses on the development and evaluation of a language model for psychological applications. However, it is important to note that multimodal LLMs, which incorporate both video and speech inputs, are potentially more effective for understanding user emotions due to their ability to process and integrate multiple modalities of data. The creation of multimodal datasets and benchmarks in the psychological domain warrants further investigation. We leave it for future work.

### v3

@@ III-D Supervised Fine-Tuning
The hyperparameters utilized for training the model are configured as follows: the learning rate is set to 1e-5, the batch size is set to 128, and the number of epochs is set to 3. We employ full parameter fine-tuning to train the model using eight A6000 GPUs, each equipped with 48GB memory.
@@ IV Benchmark
Figure 5: Examples of MCQs in the proposed benchmark, including SMCQ, MMCQ and case-based MCQs, respectively.
@@ IV-A Dataset Collection and Preprocessing
In order to accurately evaluate the model performance in the psychology domain, we introduce a benchmark that psychological professional individuals need to master. The data is sourced from publicly available examination questions. We follow a standardized data preprocessing procedure. For some data, we first need to perform optical character recognition (OCR) to convert images into text. After that, we invite several students to manually review the collected data to ensure its quality and consistency with the original document. This process involves rectifying formatting errors, eliminating duplicate questions, and rectifying any instances of garbled characters. Our proposed psychological benchmark draws inspiration from the format of the most authoritative psychological counseling examination in China, comprising two primary components: objective questions and subjective questions.
Note that the objective questions mainly comprise two types of MCQ: MCQ with only a single correct option (SMCQ) and MCQ with multiple correct options (MMCQ). The content of benchmark can be categorized into three sections: professional ethics, theoretical proficiency, and case analysis. The professional ethics section generally encompasses general principles such as honesty, integrity, professional responsibility, and social responsibility. Theoretical proficiency typically pertains to fundamental theories and concepts in psychology, including cognitive psychology, developmental psychology, and related fields. Both the professional ethics and theoretical proficiency sections contain only MCQs, designed to evaluate the mastery of knowledge within the psychological domain. For case analysis, it includes both MCQs and iterative questioning based on given case, aiming to better evaluate performance in real-world scenarios. The former focuses on assessing the grasp of professional knowledge related to the case, while the latter offers a more open-ended assessment of the model’s capability in practical utility. Finally, we obtained a total of 3863 MCQs and 20 cases with 100 QAs. An example of MCQs is shown in Fig. 5, and an example of case-based QA is presented in Fig. 6. The statistics of the benchmark are shown in Table III, detailing the distribution of questions across various categories. The questions are classified into three main categories: Level 2 Psychological Counselor Exam, Level 3 Psychological Counselor Exam, and Others. Note that, according to the official definitions, the Level 2 exam is considered more challenging than the Level 3 exam.
@@ IV-B Evaluation Metric
There are three elements in each instance of the case analysis. $\mathcal{B}$ is the description of the given case, including the background, environment, etc. of the seeker, $\mathcal{Q}$ is the question according to the case, such as the diagnosis, reason and treatment for the case and $\mathcal{A}$ is the reference answer for each questions. For QA in case analysis, we use the models to generate the response $\mathcal{O}$ according to the given $\mathcal{B}$ and $\mathcal{Q}$. We evaluate the response $\mathcal{O}$ from model with the corresponding references $\mathcal{A}$. We follow existing text generation metric, including ROUGE-1 (R-1), ROUGE-L (R-L) [61], BLEU-4 (B-4) [62], and BERTScore [63]. For the ROUGE metric, we report the $\textsc{F}_{1}$ scores for both R-1 and R-L. R-1 measures unigram overlap, which serves as an indicator of informativeness, while R-L evaluates the longest common subsequence overlap, providing an assessment of fluency. BERTScore leverages embedding from pre-trained encoder to capture the semantic and contextual information of tokens through computing cosine similarities. This metric allows for a more nuanced evaluation of text generation quality by considering the underlying meaning and context rather than just surface-level token matching.
For SMCQ, we use the standard accuracy to evaluate the model performance. For MMCQ, we employ two types of accuracy metrics: standard accuracy and elastic accuracy. Standard accuracy is defined as the ratio of completely correct answers to the total number of questions, providing a clear measure of the model’s precision by counting only fully accurate responses. Elastic accuracy, on the other hand, accounts for both fully accurate and partially correct answers, where latter contain no incorrect options but are not entirely complete. This metric offers a evaluation by recognizing the model’s ability to produce answers that are mostly correct, even if they are not fully comprehensive. The elastic accuracy for a MMCQ is computed as follows:
@@ V-A Baseline
  • BaiChuan2: This is a generation of open-source large language models launched by Baichuan Intelligent Technology, which is trained on a high-quality corpus with 2.6 trillion tokens.
@@
  • LLaMA3-Chinese [64]: This model uses LLama3 as the base model and is fine-tuned with high-quality Chinese instruction datasets with the training methods of DORA and LORA+.
  • Yi (e.g., Yi-1.5-6B, Yi-1.5-9B): These open-source large language models are trained from scratch by developers at 01.AI, which are trained with a high-quality corpus of 500B tokens.
@@
Besides, we compare with several existing psychological LLMs:
@@ V-B Evaluation Details
Prompts in our experiments usually comprise following components: a well-defined task description and a pertinent question soliciting the model’s response. For tasks involving MCQ, we extract answers from model outputs using an empirically designed regular expression.
@@ VI-B Results on Case-Based QA
We report the results of subjective questions in Table V. The results obtained from automatic evaluations using ROUGE and BLEU metrics for these models are relatively low. This may be due to the inherent limitations of these metrics in capturing the nuances and complexities of natural language generation, as they primarily focus on surface-level n-gram overlap rather than semantic equivalence and contextual appropriateness. Therefore, we also incorporate BERTScore into our evaluation framework. Most models consistently score between 63 and 65 on the BS indicator, indicating the consistency and stability of the BS evaluation. PsycoLLM demonstrates superior performance on R-1, B-4, and BERTScore, respectively and EmoLLM achieves better results on R-L. Additionally, although EmoLLM and MindChat do not perform well on MCQs, they achieve relatively strong performance in R-1, R-L, and B-4. The main reason might be that these models are fine-tuned with psychological conversation data, enabling them to better grasp case information and generate higher-quality responses. We provide a comparative example of the results from PsycoLLM and GPT-4o in Fig. 6. Both models are proficient at identifying the majority of patient issues. A notable distinction emerges in their responses: GPT-4o tends to offer more detailed and specific answers, while PsycoLLM generally provides more concise responses.
@@ VI-C Effect of Difficulty
VI-D Effect of Psychological Fine-tuning on General Ability
To evaluate the impact of fine-tuning on psychological datasets with respect to general reasoning abilities, we conducted experiments comparing PsycoLLM to its non-fine-tuned baseline model, Qwen1.5-14B. This comparison is performed across several widely used general benchmarks, including MMLU [30], CMMLU [66], GSM8K [67] and CEVAL [68]. The results are presented in Table VII. We can find that PsycoLLM and the base model exhibit comparable performance on these general tasks, with PsycoLLM achieving even better results on the GSM8K benchmark. This suggests that fine-tuning on the proposed psychological dataset does not significantly compromise general reasoning capabilities. However, we observe a slight decline in several benchmarks (e.g., MMLU, CMMLU, and CEVAL), which underscores the need to balance domain-specific fine-tuning with the preservation of general reasoning abilities. This decline may be attributed to the fact that certain tasks in these benchmarks may not been adequately considered during fine-tuning due to the differences between general and psychological domains. Additionally, there is a risk of overfitting to the psychological dataset, which could result in reduced robustness for general tasks.
VI-E Discussion
1) Model Size: On the one hand, within the same model families, models with larger parameters generally obtain better results in Table IV. For instance, models such as YI-1.5-9B and Qwen1.5-14B-Chat demonstrate better performance compared to their smaller counterparts, YI-1.5-6B and Qwen1.5-7B-Chat, respectively. This performance enhancement can be attributed to the increased capacity of models with larger parameters to capture complex patterns and effectively learn text representations from the data. This observation aligns with findings from a broad spectrum of existing studies that emphasize the advantages of larger model sizes in terms of learning capacity and performance. On the other hand, when compared models from different model families, this findings have slight difference. For example, YI-1.5-9B outperform Baichuan2-13B and Qwen1.5-14B-Chat in the MCQ evaluation. This discrepancy shows that training corpus, and other factors such as training procedures also significantly influence performance.
2) Model preference in MMCQs: In our experiments, we observe distinct response patterns from different models when addressing MMCQs. For example, GPT-3.5-turbo and GPT-4o frequently provide a single option in MMCQs, whereas Qwen-turbo tends to select all available options. Therefore, GPT-3.5-turbo and GPT-4o have a large improvements when shifting from standard accuracy measure to elastic accuracy metric. The main reason might be that, although GPT-3.5-turbo and GPT-4 are multilingual, they primarily focus on English. In English-speaking contexts, MMCQs are less prevalent compared to SMCQs. Consequently, the training data for these models may predominantly contains SMCQs and lack a substantial MMCQ dataset. This discrepancy leads the models to exhibit a tendency to select only one answer when presented with MMCQs.
3) Fine-tuned Models vs. General Models: Among these models, EmoLLM, MindChat are fine-tuned with the psychology-related dataset and others are for general domain. Specifically, EmoLLM and MindChat leverage empathetic conversation data for fine-tuning, which contributes to their relatively higher performance in Case-Based QA than general LLMs. However, due to their limited incorporation of knowledge-based data, their performance in MCQ evaluations is average.
@@ VII Conclusion
In this paper, to enhance LLM in the psychological domain, we proposed a high-quality psychological dataset. Specifically for multi-turn dialogue, we employed a comprehensive pipeline that includes stages of generation, evidence support, and refinement. For knowledge-based QA, we used a teacher-student LLM-based agent to enhance data generation. Additionally, we developed a benchmark based on psychological examinations. The proposed PsycoLLM, fine-tuned on the psychological dataset, demonstrated superior performance on the benchmark, excelling in both knowledge mastery and case analysis compared to other LLMs.
This paper primarily focuses on the development and evaluation of a language model for psychological applications. However, it is important to note that multimodal LLMs, which incorporate both video and speech inputs, are potentially more effective for understanding user emotions due to their ability to process and integrate multiple modalities of data. The creation of multimodal datasets and benchmarks in the psychological domain warrants further investigation. We leave it for future work. Besides, our current dataset generation process relies on LLMs, which may introduce biases inherent in these models. To address this challenge, it is crucial to explore more robust bias-mitigation strategies, which can further enhance the quality and reliability of the dataset, ultimately aiming to minimize the negative effects of bias associated with LLMs. This aspect will also be considered in future research.
TABLE VII: Comparison of the fine-tuned model and its base model on the general benchmark.
Model MMLU CMMLU GSM8k CEVAL
5-shot 5-shot 8-shot 5-shot
Qwen1.5-14B-Chat 66.01 75.13 68.31 76.00
PsyCoLLM 64.55 74.49 73.39 75.26